Use Gauss's approach to find the following sums (do not use a formula).
a)1+2+3+4+...101
b)1+3+5+7+...103
Thank you!

Answer by stanbon(75887)

Use Gauss's approach to find the following sums (do not use a formula).
a)1+2+3+4+...101
n = 101
S(n) = (n/2)(1st + last) = 101(102)/2 = 101*51 = 5151
-----------------------------------------------------
b)1+3+5+7+...103
n = 104/2 = 52
------
S(n) = (n/2)(1st + last) = 26(104) = 2704
============================================
Cheers,
Stan H.
===============